Circular No. 8502

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5115 SAGITTARII = NOVA SAGITTARII 2005 N. N. Samus and E. Kazarovets, Institute of Astronomy, Russian Aca-

demy of Sciences, inform us that the designation V5115 Sgr has been given to this nova (cf. IAUC 8500, 8501).

Visual magnitude estimates: Mar. 30.127 UT, 7.7 (K. Hornoch, Lelekovice, Czech Republic); 30.20, 8.0 (A. Pereira, Cabo da Roca, Portugal); 30.273, 7.7 (R. Y. Shida, São Paulo, Brazil).

SUPERNOVA 2005ay IN NGC 3938 S. Taubenberger, Max Planck Institut für Astrophysik, Garching; and

S. Benetti and A. Harutyunyan, Instituto Nazionale di Astrofisica, Osservatorio Astronomico di Padova; on behalf of the European RTN collaboration (cf. IAUC 7987), report that a spectrum of SN 2005ay (cf. IAUC 8500), obtained on Mar. 29.98 UT by A. Zurita with the Calar Alto 2.2-m telescope (+ CAFOS; range 340–880 nm), is that of a type-II supernova soon (maybe 8–10 days) after explosion. The spectrum is characterized by a blue continuum with broad P-Cyg features of the hydrogen Balmer series (the expansion velocity from Hβ is ∼ 9600 km/s). The He I 587.6-nm P-Cyg line is also present.

H. Yamaoka, Kyushu University, writes that K. Itagaki (Teppo-cho, Yamagata, Japan) found nothing at the location of SN 2005ay (cf. IAUC 8500) on a CCD image taken on Mar. 19.621 UT (limiting mag 19.0).

SATELLITES OF JUPITER Further to IAUC 8177, the IAU WGPSN has approved the following

new designations and names of satellites of Jupiter:

Jupiter XXXIX  Hegemone = S/2003 J 8
Jupiter XL  Mneme  = S/2003 J 21
Jupiter XLI  Aoede  = S/2003 J 7
Jupiter XLII  Thelxinoe = S/2003 J 22
Jupiter XLIII Arche = S/2002 J 1
Jupiter XLIV  Kallichore = S/2003 J 11
Jupiter XLV  Helike  = S/2003 J 6
Jupiter XLVI  Carpo  = S/2003 J 20
Jupiter XLVII  Eukelade = S/2003 J 1
Jupiter XLVIII  Cyllene  = S/2003 J 13

2005 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green